Find the mean, the median, and the mode for the sample data represented by the table

x 1 2 7

f 1 2 1
Answer #1

here putting above data in ascending order:

S. No x
1 1
2 2
3 2
4 7
total 12
average 3

mean =3

median =middle value =(4+1)/2 th value =2.5 th value =(2+2)/2 =2

mode =2 (as it appears maximum number of times)

> There is only one 2 in the value f x. Why you put 2 two times?

Warda Akram Sun, Jul 25, 2021 12:52 AM

> thx

gavin Mon, Aug 2, 2021 11:21 PM